OpenStudy (anonymous):

Okay so I have been sick this week and my friend brought by my homework and I don't know what to do. Help? Sole the equation" P = 2L + 2W for L

OpenStudy (anonymous):

maybe just maybe, if we subtract 2W from both sides, and then divide both sides by 2 we might get our equation

OpenStudy (anonymous):

\[P=2L+2W\] subtract 2W from both sides to get \[P-2W=2L\] what lagrange said

OpenStudy (anonymous):

then divide by 2 to get \[\frac{P-2W}{2}=L\]

OpenStudy (anonymous):

Thank you :)